OpenStudy (anonymous):

Solve the equation. (w+3)^-1/5=1/4

OpenStudy (anonymous):

Is the correct answer -3+4^1/15

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\[\frac{1}{\sqrt[5]{W+3}}=\frac{1}{4}\]

OpenStudy (anonymous):

coss time\[\sqrt[5]{W+3} =4\]\[(\sqrt{W+3})^{5}=4^{5}\] W+3=1.024 W =1024-3 W=1.021

OpenStudy (anonymous):

So my answer was wrong

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

negative exponent bring down stair change positive

OpenStudy (anonymous):

K thx

OpenStudy (anonymous):

u r we com